OpenStudy (anonymous):

the surface areas of two similar solids are 311ft^2 and 1,307ft^2. the volume of the larger solid is 1,755ft^3. what is the volume of the smaller solid?

OpenStudy (anonymous):

A) 961ft^3 B) 1755ft^3 C) 789ft^3 D) 288ft^3

OpenStudy (bahrom7893):

Okay first of all take the square root of 311ft^2 and 1307ft^2

OpenStudy (bahrom7893):

wait that won't work out.

OpenStudy (anonymous):

soo its 311 and 1307ft

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hmmmm

OpenStudy (bahrom7893):

The ratio between the volumes squared must be equal to the ratio between the areas raised to the third power

OpenStudy (anonymous):

yeah dude your talkin japaneze to me

OpenStudy (bahrom7893):

According to that: (1307/311)^3 = (1,755/x)^2

OpenStudy (bahrom7893):

74.224 = 3080025/x^2 x^2 * 74.224 = 3080025 x = 3080025/74.224

OpenStudy (bahrom7893):

wait no: x^2 = 3080025/74.224

OpenStudy (anonymous):

41496.34889

OpenStudy (anonymous):

oooh

OpenStudy (bahrom7893):

x^2 = 41 496.3489

OpenStudy (bahrom7893):

x = 203.7

OpenStudy (anonymous):

but thts not one of my answer choices

OpenStudy (bahrom7893):

yea i see that now.. Im still lookin over my work..

OpenStudy (anonymous):

its obviously not b

OpenStudy (bahrom7893):

do you know what the solid is?

OpenStudy (bahrom7893):

Okay let's try the previous method..

OpenStudy (bahrom7893):

the surface areas of two similar solids are 311ft^2 and 1,307ft^2. the volume of the larger solid is 1,755ft^3. what is the volume of the smaller solid? A) 961ft^3 B) 1755ft^3 C) 789ft^3 D) 288ft^3

OpenStudy (anonymous):

no it jus said two similar solids

OpenStudy (bahrom7893):

sqrt(311) = 17.635 sqrt(1307) = 36.152

OpenStudy (bahrom7893):

17.635/36.152 = (x/1755)^(1/3)

OpenStudy (bahrom7893):

0.116 = x/1755

OpenStudy (bahrom7893):

x = 1755*0.116 = 203.58

OpenStudy (bahrom7893):

damn it.. i got the same thing..

OpenStudy (anonymous):

can someone help me :(

OpenStudy (anonymous):

dam lol see thts wat i goy yesterday

OpenStudy (anonymous):

if you help me first

OpenStudy (anonymous):

*got

OpenStudy (bahrom7893):

jlane can u check the answer?

OpenStudy (anonymous):

how you mean?

OpenStudy (bahrom7893):

like can u check it, do u have the actual answer?

OpenStudy (anonymous):

no its a test ?

OpenStudy (bahrom7893):

oh okay..

OpenStudy (anonymous):

sorry i couldnt help you :(

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

not really bright with math :( but i would love if you guys would be kind to help me :) <3 please .

OpenStudy (bahrom7893):

You know what, let's try the last resort..

OpenStudy (bahrom7893):

maybe that's what they want us to do..

OpenStudy (bahrom7893):

Geez tabbey post that as a new thread or smth..

OpenStudy (bahrom7893):

the surface areas of two similar solids are 311ft^2 and 1,307ft^2. the volume of the larger solid is 1,755ft^3. what is the volume of the smaller solid? A) 961ft^3 B) 1755ft^3 C) 789ft^3 D) 288ft^3

OpenStudy (bahrom7893):

1755/1307 = x / 311

OpenStudy (bahrom7893):

1755*311/1307=x

OpenStudy (anonymous):

mr.ilovemath u dont have to be so fre4akin rude ohklay ! >=O i just want help just like everyone else does . :(

OpenStudy (anonymous):

@ bahrom you read my mind i feel like its a chat on here

OpenStudy (anonymous):

tabbey ur bein rude

OpenStudy (anonymous):

ho w-_- w.e man ; i didnt mean to come and start something ; u guys are just not nice .

OpenStudy (bahrom7893):

1755*311/1307 = 417.601 ... ohh man that didn't work either.. Okay guys stop arguing and Jlane can u take a picture of the test or double check the numbers..

OpenStudy (anonymous):

i been workin on this problem for two days so please

OpenStudy (anonymous):

yeah hld on

OpenStudy (bahrom7893):

I can't see..

OpenStudy (bahrom7893):

that's too small wait hold on are you doing that test from this computer?

OpenStudy (anonymous):

yeah

OpenStudy (bahrom7893):

Okay then look for a button called Print Screen. On your keyboard, it's usually in the upper right corner..

OpenStudy (anonymous):

found it

OpenStudy (bahrom7893):

then open up your test and hit the Print Screen button. That will save whatever you are seeing on your screen. Go into Paint and paste it.

OpenStudy (bahrom7893):

save the image and upload it here.

OpenStudy (anonymous):

it wont len me copy it

OpenStudy (bahrom7893):

Copy what once you hit print screen that's it! your file is copied.. then go to paint and press Ctrl+V at the same time..

OpenStudy (anonymous):

tick tock tic toc

OpenStudy (bahrom7893):

jlane where r u? why aren't u doing as I said?

OpenStudy (bahrom7893):

Just open the exam. Hit Print Screen then Open Paint, go to Edit - Paste. Save file upload it here!

OpenStudy (anonymous):

i did i got a copy an i go to post it my computer jus freexes being gay

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Hmm... better check that...

OpenStudy (bahrom7893):

Estudier please tell us how u got D

OpenStudy (anonymous):

yeah please do an it jus wnt load but the question is eye for eye and dot for dot

OpenStudy (bahrom7893):

okay.. then let's take a look at it again: the surface areas of two similar solids are 311ft^2 and 1,307ft^2. the volume of the larger solid is 1,755ft^3. what is the volume of the smaller solid? A) 961ft^3 B) 1755ft^3 C) 789ft^3 D) 288ft^3

OpenStudy (anonymous):

Actually I got 203 with 1755 * sqrt(311/1307) all cubed. Then I found http://www.algebra.com/algebra/homework/Surface-area/Surface-area.faq.question.205535.html

OpenStudy (bahrom7893):

YAY!!! Estudier u saved our BOOTIES!!

OpenStudy (bahrom7893):

the answer in the link is right!

OpenStudy (anonymous):

But the question is slightly different?

OpenStudy (bahrom7893):

No it's not!

OpenStudy (anonymous):

1037 , 1307

OpenStudy (anonymous):

So is the question mistyped?

OpenStudy (anonymous):

Bizarre. I got 203 ft^3 for that question

OpenStudy (anonymous):

Yes, that's what I posted above, I think the question is mistyped and should be 1037, not 1307.

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

my test says 1307 idk lool i hope its rite

OpenStudy (anonymous):

If it's 1037, then it's D If it's 1307, then the answer is 203, not in your list.

OpenStudy (anonymous):

Don't you know how to compute scale factor jlane?